Please be detailed thanks.
PRINTER VERSIONBACK NEXT Chapter 02, Problem 10 In reaching her destination, a backpacker walks with an averagev with an average velocity of 3.34 m/s due west, turns around, and hikes with an average velocity of 0.479 m/s due east. How far east did she walk (in kilometers)? West I East Number Units ( the tolerance is +/-596
Answer #1

We know the average velocity is defined as total traveled distance divided by total time taken, i.e

Vavg = total distance(D)/total time(t),

and from here we can write time t = D/Vavg

So, It took her t = (5530m) / (3.34m/s) = 1656 s to walk the first segment.
It took her t' to walk the second segment. For "west" negative,
-5530m + 3.34m/s*t' = -1.03m/s*(t' + 1656s)

4.37t = 3824.32

This solves to t' = 875 seconds

Then she walked east x = 0.479m/s * 875s = 419.18 m = 0.42 km

If you find this helpful, please thumbs up.